OpenStudy (anonymous):

Prove the sum of cubes of 1, ... , n equals (n*(n+1)/2)^2

OpenStudy (akshay_budhkar):

lol truthfully i just memorized the stuff for my exams nver thought of proving it in my school.. i will still try

OpenStudy (anonymous):

Here's the LaTex for the equation: \[1^{3} + 2^{3} + ... + n^{3} = (n(n+1)/2)^{2}\]

OpenStudy (anonymous):

you know you use mathematical induction Check the truth of the statement for 1: 1^3=((1*(1+1))/2)^2=1^2=1. Thus the statement is true for 1 Now, assume that for n, the statement is true And now, we prove the statement for n+1 Thus, we need to prove \[1^3+2^3+...+n^3+(n+1)^3=((n+1)(n+2)/2))^2\] Now, since \[1^3+2^3+...+n^3=(n(n+1)/2)^2\] is true then \[1^3+2^3+...+n^3+(n+1)^3=(n(n+1)/2)^2+(n+1)^3\] is true Now simplify \[(n(n+1)/2)^2+(n+1)^3\] Eventually you will get \[((n+1)(n+2)/2)^2\] Yeah, and so you can conclude that the statement is true for n+1. And so the statement has been proven

OpenStudy (akshay_budhkar):

lol this is a good method!! but we also have another method which i don't seem to remember!

OpenStudy (anonymous):

Thats mathematical induction.. pretty neat!

OpenStudy (anonymous):

Thanks Anhhuyalex! Akshay - If you find the proof without using induction, I'd certainly be interested to see it.